### v2

# AniME: Adaptive Multi-Agent Planning for Long Animation Generation

Abstract.
We present AniME, a director-oriented multi-agent system for automated long-form anime production, covering the full workflow from a story to the final video. The director agent keeps a global memory for the whole workflow, and coordinates several downstream specialized agents. By integrating customized Model Context Protocol (MCP) with downstream model instruction, the specialized agent adaptively selects control conditions for diverse sub-tasks. AniME produces cinematic animation with consistent characters and synchronized audio–visual elements, offering a scalable solution for AI-driven anime creation.
copyright: none
Figure 1. Brief Architecture of AniME
1. Introduction
Anime production is a complex, labor-intensive process encompassing multiple creative stages such as scriptwriting, storyboarding, character and scene design, animation, voice acting, and final editing. Traditional workflows require extensive manual expertise and close collaboration across diverse teams, resulting in high costs and long production cycles. Recent advances in generative AI, such as AniSora (Jiang et al., 2024) for animation generation, have demonstrated impressive capabilities in specific tasks (Du et al., 2025; Majumder et al., 2024). However, these methods each exhibit distinct strengths and weaknesses in particular domains, leading to challenges in maintaining consistency (Yang et al., 2024; Shi et al., 2025) and achieving fine-grained controllability in agent-driven video generation (Li et al., 2024; Wu et al., 2025; Xia et al., 2025). As a result, developing a fully automated system for long-form anime generation remains an open challenge, particularly in selecting appropriate control conditions and ensuring cross-stage content consistency.
In this work, we present AniME, a novel director-oriented multi-agent framework that integrates specialized agents equipped with tailored Model Context Protocol (MCP) (Anthropic, 2024) toolsets. The framework employs centralized planning and quality control to coordinate task scheduling and ensure content consistency across stages. It enables coherent long-form video generation through a combination of rich, controllable generative models and iterative feedback workflows. Furthermore, we customize the MCP protocol of downstream generative tools to explicitly annotate each tool’s domain expertise and limitations, allowing agents to select appropriate control condition models for specific sub-tasks.
Table 1. Summary of Specialized Agents and their MCP Toolsets
Agent Inputs Outputs Tools & Methods
Character Designer visual style, character description character multi-view image Text-to-image + refinement; multi-view synthesis
Scene Designer visual style, Scene description Background images, layered assets Depth-guided image generator; layout-guided image generator; relighting model
Script & Storyboard Shot description text Timeline, shot prompt, keyframes, camera motion LLM-based segmentation and tagging; camera planner; layout planner; reference image generation
Animator Keyframes/camera paths/rigs/poses/audio Frame sequences Keyframe/audio/pose/camera conditioned video generation model
Audio Production Dialogue with emotion labels, scene tags Phoneme-aligned audio speaker conditioned TTS; text/video to music model; audio mixer programs
Video Editor Frame sequences, audio stems, editorial instructions Final encoded video Transition effects; color pipeline; FFmpeg multi-pass encoding
Quality Evaluator Generated frames and assets text-to-video similarity text-to-image similarity; identity verification; AV sync checks; VLM narrative evaluator
2. Method
AniME decomposes the story-to-video task into hierarchical stages coordinated by a centralized Director Agent and executed by a set of Specialized Agents. Each agent $A_{i}$ has a well-defined input type $\mathcal{I}_{i}$, output type $\mathcal{O}_{i}$, and a local MCP toolbox $\mathcal{T}_{i}$. Communication between agents is performed via structured JSON messages. Refer the method details in the supplementary
2.1. Director Agent
The Director Agent serves as the central controller of the AniME framework, managing both the global workflow and quality assurance. It decomposes the input story into a task workflow, allocates subtasks to agents, checks the quality of each specialized agent, and maintains a global Asset Memory Bank.
1
Input: Story script $\mathcal{R}$
Output: Final video $\mathcal{V}$
2
3Initialize Director with $\mathcal{R}$, derive shots $\mathcal{H}$ and styles $\mathbf{s}_{v},\mathbf{s}_{a}$;
4 Generate task list $T$ and task dependencies graph $\mathcal{W}=(N,E)$;
5
6foreach task $t\in T$ in topological order of $\mathcal{W}$ do
7   Assign $t$ to agent $A_{i}$;
8   Agent $A_{i}$ select proper model from MCP toolset $\mathcal{T}_{i}$ according to task requirement;
9  Agent $A_{i}$ producing output $\mathcal{O}_{i}(t)$ and return to Director;
10 if Evaluated quality of $\mathcal{O}_{i}(t)$ low then
11      Request revision
12   end if
13  Store $\mathcal{O}_{i}(t)$ in Asset Memory;
14
15 end foreach
16
17Director performs final editing via Video Editor Agent;
18
19return $\mathcal{V}$;
20
ALGORITHM 1 AniME: Multi-Agent Workflow
2.1.1. Workflow of the Director
Given a long-form story $\mathcal{R}$, the Director initiates a hierarchical breakdown into scenes and shots through a segmentation process, and decides the style for visual $\mathbf{s}_{v}$ and $\mathbf{s}_{v}$ acoustic setting. Then it uses chain-of-thought prompts to generate an initial task list $T$. Each task is clearly defined with input/output specifications following the downstream agent format.
The Director maintains a workflow graph $\mathcal{W}=(N,E)$, where each node $n\in N$ corresponds to a production task (e.g., “generate character pose for shot 3”) and edges $E$ encode explicit dependencies (e.g., character design $\rightarrow$ storyboard $\rightarrow$ animation).
Table 2. Names and key fields for the Asset Memory Table
Table Name Keys
shot id, description
scene id, prompt, view_3d
character id, prompt, demo_voice, voice_prompt, 3d_view
style id, visual_style, acoustic_style
storyboard id, prompt, image_path
video id, prompt, video_path, shot_id, music_id
music id, character_id, prompt, music_path
2.1.2. Asset Memory Management
The Asset Memory Management module stores and organizes key creative assets across the production pipeline, ensuring consistency and reusability. The assets are shown in Table  2. The management is implemented using queryable database tables, enabling efficient indexing, retrieval, and update operations.
2.2. Specialized Agents and MCP Tools Selection
AniME employs a set of specialized agents designed for a key production stage with dedicated MCP toolsets. The Table 1 demonstrates a detailed overview of each agent’s inputs, outputs, and tools. Specifically, we describe the advantages and disadvantages of each tool in the MCP protocol, enabling the agent to decide on the task requirements. Please refer to the MCP Selection details in the supplementary.
The Script & Storyboard transforms narrative scripts into structured shot descriptors and visual keyframes with camera plans. The Character Designer generates canonical character assets, including multi-view portraits and identity embeddings. The Scene Designer creates backgrounds and environmental assets consistent with scene perspective and style. The Animator synthesizes animated frame sequences and performs precise lip-syncing, maintaining character identity and style coherence. The Audio Production handles expressive speech synthesis, ambient sound generation, music composition, and audio mixing. The Video Editor integrates visual and audio components, applies color grading and transitions, and produces the final encoded video. Finally, the Quality Evaluator automates multi-modal verification of visual and narrative coherence, triggering targeted revisions when necessary.
Meanwhile, animation creators can interact with the agent at any stage, providing revision suggestions and guiding the agent to produce animations that better align with the creator’s artistic style.
3. Conclusion
We presented AniME, a director-oriented multi-agent framework for automated long-form anime production. By integrating specialized agents equipped with customized MCP toolsets, the agents adaptively decide the optimistic available models and achieved consistent generation from a story to the final video.